# Dynamics of ballistic photocurrents driven by Coulomb scattering

### v1

@@ Abstract
In summary, we have observed ballistic currents generated by Coulomb scattering in rt-TDDFT simulations. These currents are comparable in magnitude to shift currents. We find that recombination currents strongly affect the total shift current, but appear to be less important for ballistic currents. These observations suggest that a multitude of mechanisms contribute to the BPVE, with consequences for understanding many experiments beyond the narrow interpretation of the excitation shift current. The ballistic current optical responsivity from Coulomb scattering measured here is similar to that from electron-phonon scattering  [24], which is expected considering that both forms of scattering have similar overall scattering rates [38], and suggesting that other forms of scattering, such as from defects or interfaces, could potentially contribute strongly to ballistic photocurrents as well. While we have shown that the Coulomb ballistic current is already present at the mean-field approximation, the effects of electron-electron correlation also warrants further study. The range of time scales revealed by these simulations, associated with the rise and changes of the various photocurrents, and transient accumulation points of carriers, suggests that ultrafast experiments will greatly aid the understanding of these phenomena.

### v2

@@ Abstract
In summary, we have observed ballistic currents generated by Coulomb scattering in rt-TDDFT simulations. These currents are comparable in magnitude to shift currents. These observations suggest that a multitude of mechanisms contribute to the BPVE, with consequences for understanding many experiments beyond the narrow interpretation of the excitation shift current. The ballistic current optical responsivity from Coulomb scattering measured here is similar to that from electron-phonon scattering  [12], which is expected considering that both forms of scattering have similar overall scattering rates [8], and suggesting that other forms of scattering, such as from defects or interfaces, could potentially contribute strongly to ballistic photocurrents as well. While we have shown that the Coulomb ballistic current is already present at the mean-field approximation, the effects of electron-electron correlation also warrants further study. Photo-Hall measurements and ultrafast time-resolved experiments will be able to distinguish the Coulomb ballistic current from other current contributions, by measuring the time scales of the rise of photocurrents, and the dynamics associated with the transient carrier accumulation points.